OpenStudy (anonymous):

Which of these recursive formulas results in this sequence? 1, 3, 6, 10, 15, ... A) an = 2a(n-1) – 1 B) an = 2a(n-1) + 1 C) an = a(n-1) – n D) an = a(n-1) + n

OpenStudy (anonymous):

HI :DDD

OpenStudy (anonymous):

hi:)

OpenStudy (tkhunny):

3-1 = 2 6-3 = 3 10-6 = 4 15-10 = 5 I'm seeing a pattern.

OpenStudy (anonymous):

I do too but I don't understand which formula would match it.

OpenStudy (tkhunny):

A) an = 2a(n-1) – 1 -- Doubles and subtracts one. Nope. B) an = 2a(n-1) + 1 -- Doubles and subtracts one. Nope. C) an = a(n-1) – n -- Subtracts one more each step. Nope. D) an = a(n-1) + n -- Adds one more each time. Maybe! Have you considered trying to regenerate the sequence?

OpenStudy (nipunmalhotra93):

the given sequence is a sequence of triangular numbers.. an=n(n+1)/2. It's better to use this than individually verifying the options.

OpenStudy (tkhunny):

Strongly Disagree. Recognizing triangular numbers works ONLY for triangular numbers. It is not a vry useful methodology. Learning what it is doing and how to build it. Learning how to use a recursive function. Understanding what you are doing. These are the more valuable methods.